A tennis ball with a mass of 60 g was moving at a velocity of 3 m/s. How fast would a ball of mass 20g have to be going to have the same momentum as the tennis
1
Expert's answer
2012-02-14T09:47:42-0500
A tennis ball with a mass of 60 g moving at a velocity of 3 m/s has a momentum

Mt = M1*V1 = 0.06 kg * 3 m/s = 0.18 kg*m/s.

So, V2 = Mt/M2 = ( 0.18 kg*m/s ) / 0.02 kg = 9 m/s.